OpenStudy (icalibear):

What is the value of the range of the function f(x) = x^2 + 2 for the domain value of 1/4?

ganeshie8 (ganeshie8):

range is just f(1/4)

OpenStudy (icalibear):

so how do i calculate this

ganeshie8 (ganeshie8):

f(x) = x^2 + 2 f(1/4) = (1/4)^2 + 2 = 1/16 + 2 =33/16